Circular No. 9043

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1213 CENTAURI G. Pojmanski, D. Szczygiel, and B. Pilecki, Warsaw University Astro-

nomical Observatory, report their discovery of a possible nova in a crowded star field on V -band CCD exposures taken with a 70-mm-aperture, 200mm-f.l. f/2.8 camera lens in the course of the All-Sky Automated Survey, providing the variable’s position as α = 13^{h}31^{m}16^{s}, δ = −63^{o}57′.6 (equinox 2000.0). ASAS-3 V -band magnitudes for the variable: May 4.168 UT, [14: (presumed; not detected); 8.235, 8.53; 11.108, 9.12; 11.140, 9.13. M. Templeton, AAVSO, reports that L. Elenin (Moscow, Russia) observed the variable at V = 9.69 on May 13.466, remotely using a 0.15-m f/7.3 refractor near Perth, W. Australia, and providing position end figures 15^{s}.76, 38′′.5 for the new object. E. Guido and G. Sostero write that they obtained unfiltered CCD images remotely with a 0.25-m f/6 telescope near Moorook, Australia, on May 13.57 that yield magnitude ∼ 8.6 for the variable and position end figures 15^{s}.77, 38′′.6; comparison with an Anglo-Australian Observatory Schmidt red plate (limiting magnitude ∼ 20), obtained on 1997 Feb. 5, shows that this position is nearly coincident with a field star with mag ∼ 15 whose position end figures are 15^{s}.68, 38′′.6 (difficult measurement due to extreme crowding of nearby field stars). Additional information was given on CBET 1800.

A. Pigulski, Astronomical Institute, Wrocław University, writes that he obtained a single spectrogram of the variable with the 1.9-m Radcliffe telescope (+ GIRAFFE échelle spectrograph; range 425–680 nm; resolution 32000) of the South African Astronomical Observatory on May 13.89 UT. The star shows very strong and broad emission in Hα, Hβ, and Hγ. The peak flux of the Hα emission is roughly ten times stronger than that of the continuum. The FWHM of the Balmer lines (as measured from Hα and Hβ) corresponds to the velocity of 2300 km/s. In addition, the Fe II emission features at 490–540 nm, characteristic for a nova at an early stage of evolution, can be clearly seen in the spectrum.

E. V. Kazarovets and N. Samus report that the GCVS team assigns the designation V1213 Cen to this nova.

COMETS 218P/LINEAR, 219P/LINEAR, AND 220P/McNAUGHT

Comet P/2009 F7 = P/2003 H4 (cf. IAUC 9038) has been assigned the permanent number 218P; comet P/2009 H1 = P/2002 LZ_{11} (cf. IAUC 9039) receives the permanent number 219P; and comet P/2009 H2 = P/2004 K2 (cf. IAUC 9040) receives the permanent number 220P.

2009 May 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green